Solve the following I VP x:]xof]

show all steps
Answer #1

X 10)= let A = Tofrnd eizen value tA TAdI) =06-A -1 4-A > (6-8)(4-d) +5 -0 2 24-6A-4ad to ts-0 > 2101 +29 =0 10 t 100-4 C29)-> (1-2) se=X2 (1-2i) u -x2 =0 >-> 5 sep + (-1+2i)H2=0 51 + (-1 +2i) (l-2u)24=0 is -- eigen vector : first solution XIt)= è (We haue XIt) = -2= X,(0) = Cie coszt t(2esinzt Cie cos 2t + Czèt sinzt %3D + C2x0. CI CUS2E %3D Xect) = é t (cos2tt2ulinzt) 5